# FairGen: Controlling Sensitive Attributes for Fair Generations in Diffusion Models via Adaptive Latent Guidance

### v1

@@ Abstract
Text-to-image diffusion models often exhibit biases toward specific demographic groups, such as generating more males than females when prompted to generate images of engineers, raising ethical concerns and limiting their adoption. In this paper, we tackle the challenge of mitigating generation bias towards any target attribute value (e.g., “male” for “gender”) in diffusion models while preserving generation quality. We propose FairGen, an adaptive latent guidance mechanism which controls the generation distribution during inference. In FairGen, a latent guidance module dynamically adjusts the diffusion process to enforce specific attributes, while a memory module tracks the generation statistics and steers latent guidance to align with the targeted fair distribution of the attribute values. Further, given the limitations of existing datasets in comprehensively assessing bias in diffusion models, we introduce a holistic bias evaluation benchmark HBE, covering diverse domains and incorporating complex prompts across various applications. Extensive evaluations on HBE and Stable Bias datasets demonstrate that FairGen outperforms existing bias mitigation approaches, achieving substantial bias reduction (e.g., $68.5\%$ gender bias reduction on Stable Diffusion 2). Ablation studies highlight FairGen’s ability to flexibly and precisely control generation distribution at any user-specified granularity, ensuring adaptive and targeted bias mitigation.
@@
$\ast$\$$\ast$\$footnotetext: Corresponding authors.$\dagger$\$$\dagger$\$footnotetext: Work done during an internship at AWS AI Labs.$\ddagger$\$$\ddagger$\$footnotetext: Work done during full-time employment at AWS AI Labs.
@@ 1 Introduction
Text-to-image diffusion models (Nichol et al., 2021; Saharia et al., 2022) have shown remarkable capabilities when generating photorealistic images from text input, leading to new real-world applications. Notably, stable diffusion models (Rombach et al., 2022; Podell et al., 2023; Esser et al., 2024) and DALL-E models (Ramesh et al., 2022; Betker et al., 2023) have gained widespread popularity, attracting millions of users and being utilized in a wide range of contexts such as reinforcement-learning based control (Pearce et al., 2023; Chi et al., 2023) and life-science (Chung et al., 2022; Cao et al., 2024).
@@
Additionally, we find that current bias evaluation benchmarks (Bakr et al., 2023; Lee et al., 2024; Cui et al., 2023; Wan and Chang, 2024; Wan et al., 2024; Luccioni et al., 2023; Naik and Nushi, 2023) exhibit three major limitations: a narrow range of domains, overly simplistic input prompt structures, and a limited set of attributes. To address these shortcomings, we propose a holistic bias evaluation benchmark HBE (in Section 3) that encompasses a wider array of domains, prompt structures, and sensitive attributes compared to previous benchmarks. Our experiments reveal that while state of the art bias-mitigation approaches excel in widely used bias evaluation benchmarks (e.g., Stable Bias Luccioni et al. (2023)), their performance drop significantly in HBE, proving the rigor of the HBE dataset (Section 5.1).
We evaluate FairGen against several state of the art baselines on the HBE and Stable Bias datasets and find that FairGen outperforms all baselines in both datasets in bias reduction and quality preservation. In summary, our major contributions and findings are as follows – (a) We define the novel problem of generating images by adhering to a target fair distribution of attributes. (b) We propose FairGen, a novel inference time approach for generating high quality images by adhering to the target distribution of attributes. (c) We propose HBE, a novel and comprehensive benchmark for assessing bias in diffusion models. (d) Extensive experimental evaluations show that FairGen outperforms SOTA bias-mitigation methods in terms of bias reduction and demonstrates greater effectiveness in scenarios involving the interplay of multiple attributes (Table 2). (e) FairGen provides an adaptable mechanism for controlling generation distributions at different target distribution levels compared to SOTA methods (Tables 3).
@@ 2 FairGen
We first introduce our fair diffusion model generation pipeline FairGen in Section 2.1, which consists of a latent guidance module and an indicator guidance module. In Section 2.2, we describe the functionality and training process of the latent guidance module, which generates adaptive guidance for specific attributes in the latent space. Section 2.3 details the indicator guidance module, which produces scalar guidance directions to enforce attribute values and achieve the target generation distribution.
@@ 2.1 Overview of FairGen
In this paper, we study the problem of generating high quality images by preserving a target distribution of different attributes present in the image (e.g., generating images of males and females with equal probability with a particular occupation). Existing bias mitigation methods using prompt intervention tend to degrade generation quality due to modification of the input prompts (Bakr et al., 2023; Lee et al., 2024; Cui et al., 2023; Wan and Chang, 2024; Wan et al., 2024; Luccioni et al., 2023; Naik and Nushi, 2023) and finetuning-based methods (Orgad et al., 2023; Shen et al., 2023; Zhang et al., 2023) degrade image quality due to fitting to subdomains. We experimentally verify this phenomenon in such approaches in Section 5.1. Moreover, finetuning based approaches require additional training to adapt to different target distribution of attributes. Therefore, we propose FairGen to impose fair generations via guidance in diffusion latent space and to flexibly control the target generation distributions at inference time.
@@
Formally, at diffusion sampling step $t$, the diffusion noise direction $\bm{\epsilon}_{\theta}(\bm{x}_{t},\bm{c})$ is given by a noise estimation network $\epsilon_{\theta}$, parameterized by $\theta$, and conditioned on the latent state $\bm{x}_{t}$ at step $t$ and the input prompt $\bm{c}$. The attribute guidance direction consists of two components: (1) a scalar guidance direction $I(\bm{c},\mathcal{M},(a_{1},a_{2}))\in\{-1,1\}$, which depends on the input prompt $\bm{c}$, an auxiliary memory module $\mathcal{M}$ containing generation statistics, and the potential attribute values for manipulation $a_{1},a_{2}$ (assuming binary value attribute for brevity here); and (2) an adaptive latent guidance direction $f_{\text{ALD}}(\bm{x}_{t},\bm{c},(a_{1},a_{2}))$, produced by a trained guidance network $f_{\text{ALD}}$, which depends on the latent state $\bm{x}_{t}$, the input prompt $\bm{c}$, and the potential attributes. The final attribute-aware noise direction is defined as follows.
@@
$+(1-\gamma)\underbrace{I(\bm{c},\mathcal{M},(a_{1},a_{2}))}_{\text{Scalor Guidance Direction}}\cdot\underbrace{f_{\text{ALD}}(\bm{x}_{t},\bm{c},(a_{1},a_{2}))}_{\text{Adaptive Latent Guidance Direction}}$
@@
The scalar guidance direction $I(\bm{c},\mathcal{M},(a_{1},a_{2}))$ acts as an indicator guidance model that determines the scalar for the guidance direction (e.g., assigning $1$ for male guidance and $-1$ for female guidance) based on the memory module $\mathcal{M}$. The adaptive latent guidance direction $f_{\text{ALD}}(\bm{x}_{t},\bm{c},(a_{1},a_{2}))$ provides the noise estimate required to modify attribute value $a_{1}$ towards $a_{2}$, conditioned on the latent variable $\bm{x}_{t}$ and the prompt $\bm{c}$.
This formulation extends to multiple multi-dimensional attributes as follows.
@@
Here, $\bm{\mathcal{A}}$ represents a set of multi-dimensional attributes (e.g., gender, race, age), and $a_{i}$ and $a_{j}$ are attribute values within the attribute $\mathcal{A}$. Figure 1 shows the overview of the proposed method.
@@ 2.2 Adaptive Latent Guidance Module
This paradigm ensures that the attribute-aware guidance prompts $K(\bm{c},a_{i})$ and $K(\bm{c},a_{j})$ share similar patterns while differing only in their target attributes. As a result, the corresponding noise predictions $\bm{\epsilon}_{\theta}(\bm{x}_{t},K(\bm{c},a_{i}))$ and $\bm{\epsilon}_{\theta}(\bm{x}_{t},K(\bm{c},a_{j}))$ reside in the same space, and their difference aligns orthogonally to the diffusion noise estimate direction $\bm{\epsilon}_{\theta}(\bm{x}_{t},\bm{c})$. This idea is inspired by findings in multi-task learning, where enforcing orthogonality between task-specific directions has been shown to enhance generalization and improve task adaptation Wang et al. (2023). We fine-tune an LLM as the attribute-aware generator $L$ in two steps.
@@
(2) Direct Policy Optimization (DPO): Since SFT only enables the model to generate format-correct guidance prompts, it may not be effective for diffusion model guidance in practice. Hence, we further refine $L$ using a DPO step Rafailov et al. (2024). We first generate a collection of output guidance prompts using the SFT endpoint of $L$ and evaluate them on the validation split of HBE dataset introduced in Section 3. Each candidate guidance prompt sampled from SFT endpoint is assigned a utility score which is a convex combination of bias score and image quality score on a surrogate model, Stable Diffusion 2 (details in Section 4). Using these rewards, we label the outputs based on the 50% quantile, distinguishing positive and negative samples and use the positive-negative pairs for performing DPO on $L$. This enables the generator to discern nuanced differences among format-correct guidance prompts, thereby enhancing attribute control and image quality.
@@ 2.3 Indicator Guidance Module
In this section, we describe how FairGen generates the scalar guidance direction $I(\bm{c},\mathcal{M},(a_{i},a_{j}))\in\{+1,-1\}$, which determines the target attribute value to enforce (i.e., $a_{i}$ or $a_{j}$). Specifically, it dictates the direction of the current generation, where $+1$ steers towards attribute value $a_{i}$ and $-1$ towards attribute value $a_{j}$. This decision process is adaptively influenced by the input text prompt $\bm{c}$ and the memory $\mathcal{M}$, which maintains generation statistics.
@@ 3 Holistic Bias Evaluation Benchmark
To fairly evaluate bias in diffusion models, it is essential to ensure that the benchmark is comprehensive and aligns with real-world scenarios. However, existing bias evaluation benchmarks suffer from three major limitations – (1) They predominantly focus on a narrow range of domains by ignoring many crucial ones. For instance, benchmarks like HRS (Bakr et al., 2023) and PST (Wan and Chang, 2024) primarily assess occupation-based biases, however, do not study some crucial domains such as healthcare, finance, and daily activities. (2) They rely on overly simplistic input prompt structures (e.g., “Photo portrait of a <objective>”), failing to capture the complexity of real-world user inputs, which often involve nuanced and context-rich descriptions. Benchmarks such as HEIM (Lee et al., 2024) and StableBias (Luccioni et al., 2023) focus predominantly on basic phrases, offering little challenge in interpreting prompts with more intricate, scenario-based descriptions. (3) Many benchmarks (Wan and Chang, 2024; Lee et al., 2024) consider only a limited set of sensitive attributes, focusing primarily on gender and race while neglecting other crucial attributes such as age. These limitations raise concerns that diffusion models deemed fair by existing benchmarks may still produce unintended biases and stereotypes when deployed in diverse real-world scenarios and user queries.
@@
To address these shortcomings, we introduce the “Holistic Bias Evaluation Benchmark” (HBE). HBE expands the scope of domains, prompt structures, and sensitive attributes beyond existing benchmarks. Specifically, we develop a set of $2000$ prompts covering diverse domains, including occupations, education, healthcare, criminal justice, finance, politics, technology, sports, daily activities, and personality traits. Notably, HBE incorporates underexplored domains such as criminal justice, technology, and finance, ensuring a more holistic assessment of bias across societal structures. Additionally, HBE features complex prompt structures, including scenario-based descriptions, which provide a more rigorous evaluation compared to static prompts that merely describe individuals (examples in Appendix A). Hence, unlike prior benchmarks that rely primarily on simplistic prompts, HBE integrates both simple and complex input structures to better reflect real-world user interactions.
@@
We construct the dataset through the following steps – (1) We use the Mistral-7B-Instruct-v0.2 model to identify key objectives within different domains (e.g., various diseases in healthcare or political positions in the politics domain). (2) The same model is then used to generate scenario-based prompts incorporating these objectives. (3) We conduct careful human checks to ensure the prompts are of high-quality and diverse. (4) Finally, we partition the 2,000 prompts into training (40%), validation (10%), and test (50%) sets. We provide prompt structure for the above process in Appendix D.
@@
To highlight the advantages of HBE over existing benchmarks, we provide a comparative analysis in Table 1, showcasing its broader domain coverage and richer prompt structures. We present examples from the HBE benchmark in Appendix A.
@@ 4 Experimental Setting
Evaluation Metrics. We use the bias score ($B$) to assess the generation bias of diffusion models and the quality score ($Q$) to evaluate the visual quality of generated images. The bias score $B$ quantifies the absolute difference between the actual and target proportions of a specific group in the generated images (e.g., the proportion of generated images of males versus the target proportion). The quality score $Q$ measures how well the generated images align with the user input prompt. Specifically, we compute $Q$ using the CLIP score between the generated images and the corresponding prompt, following Luccioni et al. (2023). More formally, we define the text-to-image model as a mapping $M:{\mathcal{V}}\to{\mathcal{Y}}$, where ${\mathcal{V}}$ represents the input text space and ${\mathcal{Y}}$ denotes the generated image space. Let $\mathcal{A}$ be the set of all possible values for a sensitive attribute (e.g., $\mathcal{A}=\{\text{male},\text{female}\}$ for gender). We denote the test set of $N$ input prompts as $\{v_{n}\}_{n=1}^{N}$, where $v_{n}\in{\mathcal{V}}$. A discriminator $D:{\mathcal{Y}}\to{\mathcal{A}}$ is used to identify the sensitive attributes in the generated images. The bias score $B$ is then defined as:
@@
Here, $P_{t}$ is the target proportion of attribute $a_{i}$. The probability ${\mathbb{P}}[\cdot]$ is estimated by Monte-Carlo methods with $T$ times of sampling ($T=10$ across the evaluations). In the multi-attribute controlling case, we further take the expectation over the set of sensitive attributes that we want to control.
Training the Attribute-Aware Generator $L$. We use the training and validation sets of the HBE dataset to train Mistral-7B-Instruct-v0.2111https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 as $L$. We use LoRA (Hu et al., 2021) during SFT and DPO.
Dataset and Models. We evaluate FairGen and other bias mitigation baselines on HBE and the Stable Bias (Luccioni et al., 2023) datasets. We consider two text-to-image diffusion models: stable diffusion 2 (SD2) (Rombach et al., 2022) and stable diffusion XL (SDXL) (Podell et al., 2023). We implement the attribute discrimination model $D(\cdot)$ following (Luccioni et al., 2023; Bakr et al., 2023), where the attributes are discriminated by question-answering using the vision-language model InstructBLIP-$2$https://huggingface.co/nnpy/Instruct-blip-v2. We validate the efficacy of $D(\cdot)$ through human evaluation (Appendix C).
@@
Table 3: Bias scores $B$ ($\downarrow$) for different target proportion $P_{t}$ of attribute male on HBE benchmark. The average (Avg) and standard deviation (Std) of the bias scores are reported in the last two columns.

### v2

@@ Abstract
Text-to-image diffusion models often exhibit biases toward specific demographic groups, such as generating more males than females when prompted to generate images of engineers, raising ethical concerns and limiting their adoption. In this paper, we tackle the challenge of mitigating generation bias towards any target attribute value (e.g., “male” for “gender”) in diffusion models while preserving generation quality. We propose FairGen, an adaptive latent guidance mechanism which controls the generation distribution during inference. In FairGen, a latent guidance module dynamically adjusts the diffusion process to enforce specific attributes, while a memory module tracks the generation statistics and steers latent guidance to align with the targeted fair distribution of the attribute values. Furthermore, we address the limitations of existing datasets by introducing the Holistic Bias Evaluation (HBE) benchmark, which covers diverse domains and incorporates complex prompts to assess bias more comprehensively. Extensive evaluations on HBE and Stable Bias datasets demonstrate that FairGen outperforms existing bias mitigation approaches, achieving substantial bias reduction (e.g., $68.5\%$ gender bias reduction on Stable Diffusion 2). Ablation studies highlight FairGen’s ability to flexibly control the output distribution at any user-specified granularity, ensuring adaptive and targeted bias mitigation.
@@
[…]
@@ 1 Introduction
Text-to-image diffusion models (Nichol et al., 2021; Saharia et al., 2022) have shown remarkable capabilities when generating photorealistic images from text input, leading to new real-world applications. Notably, stable diffusion models (Rombach et al., 2022; Podell et al., 2023; Esser et al., 2024a) and DALL-E models (Ramesh et al., 2022; Betker et al., 2023) have gained widespread popularity, attracting millions of users and being utilized in a wide range of contexts such as reinforcement-learning based control (Pearce et al., 2023; Chi et al., 2023) and life-science (Chung et al., 2022; Cao et al., 2024).
@@
Additionally, we find that current bias evaluation benchmarks (Bakr et al., 2023; Lee et al., 2024; Cui et al., 2023; Wan and Chang, 2024; Wan et al., 2024; Luccioni et al., 2023; Naik and Nushi, 2023) exhibit three major limitations: a narrow range of domains, overly simplistic input prompt structures, and a limited set of attributes. To address these shortcomings, we propose a holistic bias evaluation benchmark HBE in Section 3 that encompasses a wider array of domains, prompt structures, and sensitive attributes compared to previous benchmarks. Our experiments reveal that while state of the art bias-mitigation approaches excel in widely used bias evaluation benchmarks (e.g., Stable Bias Luccioni et al. (2023)), their performance drops significantly in HBE, proving the rigor of the HBE dataset (Section˜5.1).
We evaluate FairGen against several state of the art baselines on the HBE and Stable Bias datasets and find that FairGen outperforms all baselines in both datasets in bias reduction and quality preservation. In summary, our major contributions and findings are as follows – (a) We define the novel problem of generating images by adhering to a target fair distribution of attributes. (b) We propose FairGen, a novel inference time approach for generating high quality images by adhering to the target distribution of attributes. (c) We propose HBE, a novel and comprehensive benchmark for assessing bias in diffusion models. (d) Extensive experimental evaluations show that FairGen outperforms SOTA bias-mitigation methods in terms of bias reduction and demonstrates greater effectiveness in scenarios involving the interplay of multiple attributes (Table˜2). (e) FairGen provides an adaptable mechanism for controlling generation distributions at different target distribution levels compared to SOTA methods (Tables 3).
@@ 2 FairGen
We first introduce our fair diffusion model generation pipeline FairGen in Section˜2.1, which consists of a latent guidance module and an indicator guidance module. In Section˜2.2, we describe the functionality and training process of the latent guidance module, which generates adaptive guidance for specific attributes in the latent space. Section˜2.3 details the indicator guidance module, which produces scalar guidance directions to enforce attribute values and achieve the target generation distribution.
@@ 2.1 Overview of FairGen
In this paper, we study the problem of generating high-quality images by preserving a target distribution of different attributes present in the image (e.g., generating images of males and females with equal probability with a particular occupation). Existing bias mitigation methods using prompt intervention tend to degrade generation quality due to modification of the input prompts (Bakr et al., 2023; Lee et al., 2024; Cui et al., 2023; Wan and Chang, 2024; Wan et al., 2024; Luccioni et al., 2023; Naik and Nushi, 2023) and finetuning-based methods (Orgad et al., 2023; Shen et al., 2023; Zhang et al., 2023) degrade image quality due to fitting to subdomains. We experimentally verify this phenomenon in such approaches in Section 5.1. Moreover, finetuning based approaches require additional training to adapt to different target distribution of attributes. Therefore, we propose FairGen to impose fair generations via guidance in diffusion latent space and to flexibly control the target generation distributions at inference time.
@@
Formally, at diffusion sampling step $t$, the diffusion noise direction $\bm{\epsilon}_{\theta}(\bm{x}_{t},\bm{c})$ is given by a noise estimation network $\epsilon_{\theta}$, parameterized by $\theta$, and conditioned on the latent state $\bm{x}_{t}$ at step $t$ and the input prompt $\bm{c}$. The attribute guidance direction consists of two components: (1) a scalar guidance direction $I(\bm{c},\mathcal{M},(a_{1},a_{2}))\in\{-1,1\}$, which depends on the input prompt $\bm{c}$, an auxiliary memory module $\mathcal{M}$ containing generation statistics, and the potential attribute values for manipulation $a_{1},a_{2}$ (assuming binary value attribute for brevity here); and (2) an adaptive latent guidance direction $f_{\text{ALD}}(\bm{x}_{t},\bm{c},(a_{1},a_{2}))$, produced by a trained guidance network $f_{\text{ALD}}$, which depends on the latent state $\bm{x}_{t}$, the input prompt $\bm{c}$, and the specified attributes. The final attribute-aware noise direction is defined as follows:
@@
$+(1-\gamma)\underbrace{I(\bm{c},\mathcal{M},(a_{1},a_{2}))}_{\text{Scalar Guidance Direction}}\cdot\underbrace{f_{\text{ALD}}(\bm{x}_{t},\bm{c},(a_{1},a_{2}))}_{\text{Adaptive Latent Guidance Direction}}$
@@
The scalar guidance direction $I(\bm{c},\mathcal{M},(a_{1},a_{2}))$ acts as an indicator guidance model that determines the scalar for the guidance direction (e.g., assigning $1$ for male guidance and 1-1 for female guidance) based on the memory module $\mathcal{M}$. The adaptive latent guidance direction $f_{\text{ALD}}(\bm{x}_{t},\bm{c},(a_{1},a_{2}))$ provides the noise estimate required to modify attribute value $a_{1}$ towards $a_{2}$, conditioned on the latent variable $\bm{x}_{t}$ and the prompt $\bm{c}$.
This formulation extends to multiple multi-dimensional attributes as follows:
@@
Here, $\bm{\mathcal{A}}$ represents a set of multi-dimensional attributes (e.g., gender, race, age), and $a_{i}$ and $a_{j}$ are attribute values within the attribute $\mathcal{A}$. Figure˜1 shows the overview of the proposed method.
@@ 2.2 Adaptive Latent Guidance Module
This paradigm ensures that the attribute-aware guidance prompts $K(\bm{c},a_{i})$ and $K(\bm{c},a_{j})$ share similar patterns while differing only in their target attributes. As a result, the corresponding noise predictions $\bm{\epsilon}_{\theta}(\bm{x}_{t},K(\bm{c},a_{i}))$ and $\bm{\epsilon}_{\theta}(\bm{x}_{t},K(\bm{c},a_{j}))$ reside in the same space, and their difference is orthogonal to the diffusion noise estimate direction $\bm{\epsilon}_{\theta}(\bm{x}_{t},\bm{c})$. This idea is inspired by findings in multi-task learning, where enforcing orthogonality between task-specific directions has been shown to enhance generalization and improve task adaptation Wang et al. (2023). We fine-tune an LLM as the attribute-aware generator $L$ in two steps.
@@
(2) Direct Policy Optimization (DPO): Since SFT only enables the model to generate format-correct guidance prompts, it may not be effective for diffusion model guidance in practice. Hence, we further refine $L$ using a DPO step Rafailov et al. (2024). We first generate a collection of output guidance prompts using the SFT endpoint of $L$ and evaluate them on the validation split of HBE dataset introduced in Section˜3. Each candidate guidance prompt sampled from SFT endpoint is assigned a utility score which is a convex combination of bias score and image quality score on a surrogate model, Stable Diffusion 2 (details in Section˜4). Using these rewards, we label the outputs based on the 50% quantile, distinguishing positive and negative samples and use the positive-negative pairs for performing DPO on $L$. This enables the generator to discern nuanced differences among format-correct guidance prompts, thereby enhancing attribute control and image quality.
@@ 2.3 Indicator Guidance Module
In this section, we describe how FairGen generates the scalar guidance direction $I(\bm{c},\mathcal{M},(a_{i},a_{j}))\in\{+1,-1\}$, which determines the target attribute value to enforce (i.e., $a_{i}$ or $a_{j}$). Specifically, it dictates the direction of the current generation, where +1+1 steers towards attribute value $a_{i}$ and 1-1 towards attribute value $a_{j}$. This decision process is adaptively influenced by the input text prompt $\bm{c}$ and the memory $\mathcal{M}$, which maintains generation statistics.
@@ 3 Holistic Bias Evaluation Benchmark
To fairly evaluate bias in diffusion models, it is essential to ensure that the benchmark is comprehensive and aligns with real-world scenarios. However, existing bias evaluation benchmarks suffer from three major limitations – (1) They predominantly focus on a narrow range of domains by ignoring many crucial ones. For instance, benchmarks like HRS (Bakr et al., 2023) and PST (Wan and Chang, 2024) primarily assess occupation-based biases, however, overlook some crucial domains such as healthcare, finance, and daily activities. (2) They rely on overly simplistic input prompt structures (e.g., “Photo portrait of a <objective>”), failing to capture the complexity of real-world user inputs, which often involve nuanced and context-rich descriptions. Benchmarks such as HEIM (Lee et al., 2024) and StableBias (Luccioni et al., 2023) focus predominantly on basic phrases, offering little challenge in interpreting prompts with more intricate, scenario-based descriptions. (3) Many benchmarks (Wan and Chang, 2024; Lee et al., 2024) consider only a limited set of sensitive attributes, focusing primarily on gender and race while neglecting other crucial attributes such as age. These limitations raise concerns that diffusion models deemed fair by existing benchmarks may still produce unintended biases when deployed in diverse, real-world scenarios.
@@
To address these shortcomings, we introduce the “Holistic Bias Evaluation Benchmark” (HBE). HBE expands the scope of domains, prompt structures, and sensitive attributes beyond existing benchmarks. Specifically, we develop a set of $2000$ prompts covering diverse domains, including occupations, education, healthcare, criminal justice, finance, politics, technology, sports, daily activities, and personality traits. Notably, HBE incorporates underexplored domains such as criminal justice, technology, and finance, ensuring a more holistic assessment of bias across societal structures. Additionally, HBE features complex prompt structures, including scenario-based descriptions, which provide a more rigorous evaluation compared to static prompts that merely describe individuals (shown by examples in Appendix˜A). Hence, unlike prior benchmarks that rely primarily on simplistic prompts, HBE integrates both simple and complex input structures to better reflect real-world user interactions.
@@
We construct the dataset through the following steps – (1) We use the Mistral-7B-Instruct-v0.2 model to identify key objectives within different domains (e.g., various diseases in healthcare or political positions in the politics domain). (2) The same model is then used to generate scenario-based prompts incorporating these objectives. (3) We conduct careful human checks to ensure the prompts are of high-quality and diverse. (4) Finally, we partition the 2,000 prompts into training (40%), validation (10%), and test (50%) sets. We provide prompt structure for the above process in Appendix˜D.
@@
To highlight the advantages of HBE over existing benchmarks, we provide a comparative analysis in Table˜1, showcasing its broader domain coverage and richer prompt structures. We present examples from the HBE benchmark in Appendix˜A.
@@ 4 Experimental Setting
Evaluation Metrics. We use the bias score ($B$) to assess the generation bias of diffusion models and the quality score ($Q$) to evaluate the visual quality of generated images. The bias score $B$ quantifies the absolute difference between the actual and target proportions of a specific group in the generated images (e.g., the proportion of generated images of males versus the target proportion). The quality score $Q$ measures how well the generated images correspond to the user input prompt. Specifically, we compute $Q$ using the CLIP score between the generated images and the corresponding prompt, following Luccioni et al. (2023). More formally, we define the text-to-image model as a mapping $M:{\mathcal{V}}\to{\mathcal{Y}}$, where $\mathcal{V}$} represents the input text space and $\mathcal{Y}$} denotes the generated image space. Let $\mathcal{A}$ be the set of all possible values for a sensitive attribute (e.g., $\mathcal{A}=\{\text{male},\text{female}\}$ for gender). We denote the test set of $N$ input prompts as $\{v_{n}\}_{n=1}^{N}$, where $v_{n}\in{\mathcal{V}}$. A discriminator $D:{\mathcal{Y}}\to{\mathcal{A}}$ is used to identify the sensitive attributes in the generated images. The bias score $B$ is then defined as:
@@
Here, $P_{t}$ is the target proportion of attribute $a_{i}$. The probability $\mathbb{P}$$\cdot$] is estimated by Monte-Carlo methods with $T$ times of sampling ($T=10$ across the evaluations). In the multi-attribute controlling case, we further take the expectation over the set of sensitive attributes that we want to control.
[…]
@@
SD-3.5-large Vanilla generation 0.653 0.358 0.536 0.395 0.734 0.357 0.732 0.387
Prompt intervention 0.602 0.336 0.482 0.363 0.650 0.326 0.554 0.342
Finetune-based 0.402 0.312 0.332 0.352 0.552 0.327 0.454 0.350
FairDiffusion 0.583 0.325 0.387 0.362 0.536 0.345 0.532 0.357
FairGen (ours) 0.118 0.346 0.194 0.398 0.381 0.346 0.397 0.385
Table 3: Bias scores $B$ ($\downarrow$) for different target proportion $P_{t}$ of attribute male on HBE benchmark with SD2. The average (Avg) and standard deviation (Std) of the bias scores are reported in the last two columns.